How many gallons of 90% antifreeze would i add to 70 gallons of 20% antifreeze to get a 80% mixture
Answer by robertb(5830)

0.20*70 + 0.90*x = 0.80*(70 + x)
==> 14 + 0.90x = 56 + 0.80x
==> 0.10x = 42
==> x = 420 gallons of 90% antifreeze must be added.